If the length of simple pendulum is increased by 44% then what is the change in the time period of the pendulum?
(a) 22 % (b) 20 %
(c) 33 % (d) 44 %

Grade:Upto college level

3 Answers

vaibhav patil
3 years ago

The equation of line joining (x1,y1) and (x2,y2) is yy1y2y1=xx1x2x1

Hence equation of line joining (1,3) and (2,7) is

y373=x121 or y34=x11

i.e. 4x4=y3 or y=4x1

Solution of equations 3x+y=9 and y=4x1 gives point of intersection. Putting second equation in first we get

3x+4x1=9 or x=107 and y=4×1071=337

i.e. point of intersection is (107,337)

Now distance of (107,337) and (1,3) is

(1071)2+(3373)2=949+14449=1537

and distance of (107,337) and (2,7) is

(1072)2+(3377)2=1649+25649=2727

and ratio is 153272=17×3×317×4×4=34

Hence, the line joining the points (1,3) and (2,7) is divided by the line 3x+y=9 in the ratio of 3:4.

The equation of line joining (x1,y1) and (x2,y2) is yy1y2y1=xx1x2x1

Hence equation of line joining (1,3) and (2,7) is

y373=x121 or y34=x11

i.e. 4x4=y3 or y=4x1

Solution of equations 3x+y=9 and y=4x1 gives point of intersection. Putting second equation in first we get

3x+4x1=9 or x=107 and y=4×1071=337

i.e. point of intersection is (107,337)

Now distance of (107,337) and (1,3) is

(1071)2+(3373)2=949+14449=1537

and distance of (107,337) and (2,7) is

(1072)2+(3377)2=1649+25649=2727

and ratio is 153272=17×3×317×4×4=34

Hence, the line joining the points (1,3) and (2,7) is divided by the line 3x+y=9 in the ratio of 3:4.

vaibhav patil
3 years ago
t1=2pisqartl/g
 
t2=1.2(2pisqartl/g)
then the diffrence is 20% 
...............................................................................................................................................................................................
Vikas TU
3 years ago
Let intial length of pendulum is L
Then, time period , T = 2pi * sqrt(L/g) 
Here g is acceleration due to gravity .
Now, length of pendulum is increased by 44%
∴New length of pendulum = L + L × 44/100 = 1.44L
Now, new time period , 
T' = 1.2T 
∴ % change in time period = 
= (1.2T - T)/T × 100
= 0.2 × 100
= 20%
Hence, option (b) is correct